# Explicit core-hole single-particle methods for - and -edge X-ray absorption and electron energy-loss spectra

### v1

@@ Abstract
Single-particle methods based on Kohn-Sham unoccupied states to describe near-edge X-ray absorption (XAS) spectra are routinely applied for the description of -edge spectra, as there is no complication due to spin-orbit (SO) coupling. - and -edge spectra are often addressed via variants of time-dependent density functional theory (TDDFT) based on SO calculations. Here, we present a computationally efficient implementation based on single-particle calculations with core holes within the frozen-core approximation. Combined with a semiempirical energy shift and a fixed spin-orbit splitting, this allows for a prediction of experimental spectra on the absolute energy scale. Such spectra are compared to linear-response TDDFT for molecules and show similar or even better match with experiment, except for multiplet effects that are not covered by the single-particle approximation. A similar picture emerges for solids, where good qualitative and sometimes even quantitative agreement to XAS and electron energy-loss spectra is achieved.
@@ I Introduction
There are several possible approaches for the simulation of XAS spectra as summarized in comprehensive reviews from different viewpoints[14, 15, 16, 17, 4]. The most accurate calculations are based on equation of motion-coupled cluster theory (EOM-CCSD), providing reliable spectra both in terms of shape[18, 19] as well as excitation energies[20, 21]. However, such approaches are only applicable to smaller molecules due to their extreme computational cost.
The workhorse for describing excitations in larger systems is density functional theory (DFT). Here, the involvement of (highly) excited states suggests the use of time-dependent DFT (TDDFT)[22, 23]. However, TDDFT is also computationally expensive, and it has limitations in particular for the description of solids. An alternative to TDDFT are single-particle core-hole methods[17, 24, 25] where the unoccupied states in the field of a core hole are interpreted as excitations. These methods are usually applied to -edge spectra[25, 26, 27, 13, 17, 27] due to the difficulties of dealing with the angular momentum of the initial state and with spin-orbit (SO) coupling[17], although several -edge EELS simulations using the single-particle density of states[28] have notably been reported[29, 30, 31, 11].
@@
In this article, we apply single-particle methods to simulate - and -edge NEXAFS spectra, specifically focusing on 2 and 3 transitions. We compare our approach to experiment for molecules as well as for solids in order to explore their strengths and weaknesses, and also compare to TDDFT spectra in the case of molecules. The manuscript first discusses the picture underlying the description of NEXAFS spectra with an emphasis on comparing the transition from the many-particle to the single-particle viewpoint. A simple treatment of SO coupling is also presented. Then the predictions of the methods are compared to experiment both for molecules as well as for solid systems. After showing that the our methodology works extremely well also for single-atom EEL spectra, we finish with conclusions.
@@ II X-ray absorption
The process of X-ray absorption can be interpreted in different approximations, as depicted schematically in Fig. 1. Generally, an X-ray photon is completely absorbed by the molecule or solid under observation, leading to a highly excited state of the system due to the large photon energy. The absorption cross-sections are proportional to the unitless oscillator strength[24, 35, 36, 37]
where is the electron mass, the unit charge, the reduced Planck constant and are Cartesian coordinates. is the energy of the ground state, the energy of the core-excited state of index (there are many excited states forming the full XAS-spectrum) and denote their corresponding many-particle wave functions. The dipole operator in length form contains the contributions of all electrons. The relevant coordinate is chosen by the polarization of the X-ray photon . A specific appears in angle-dependent X-ray spectroscopy[38] or X-ray circular dichroism[16]. The target molecules or solids are usually in random orientations, such that the averaged intensity, i.e. is measured.
The calculation of transition energies and oscillator strengths in Eq. (1) can be performed in various approximations, where the most exact but also most demanding is the use of many-body wave functions as provided by coupled cluster (CC) descriptions for ground and excited states. The equivalent could be achieved by TDDFT in case the exact (frequency dependent[39]) exchange-correlation kernel would be known. Practical TDDFT calculations, in particular based on gradient-corrected kernels as applied here, do not, however, necessarily lead to accurate transition energies[40].
XAS is often interpreted within the single-particle picture as transitions of a core electron to an empty orbital (or the conduction band for solids). The oscillator strength reduces in this picture to[41]
@@
These single-particle energies may be again evaluated in different approximations that mainly differ by the occupation of the core state as well as by the occupation of the lowest unoccupied state (LUMO). These choices not only influence the energy values, but also change the single-particle states in Eq. (2), where we have suppressed the labels for clarity. More details about the implementation can be found in the Supplementary Information (SI) Sec. II.
Regardless of whether single-particle or many-particle pictures are adopted, the calculated spectra are usually rigidly shifted to match their lowest energy peak with that of the experiment[42]. With the exception of computationally expensive EOM-CCSD calculations[22] these shifts often amount to several eV.
@@
that can be used to correct the raw calculated energies as indicated in Fig. 1. This method has successfully predicted the XPS spectra of ionic clusters [46] and carbon materials[47, 48], as well as to S 1 NEXAFS spectra [13, 3]. We will see below that there is also substantial uncertainty in experimental absolute energies. The shift can therefore be seen as an average over experimental knowledge. There are hints that at least a large part of the shift stems from the use of the frozen-core approximation, as all-electron calculations seem to have negligible shifts at least for light elements[49, 50].
@@
Having obtained an accurate value for the first transition energy, i.e. to the LUMO in NEXAFS spectra, the question appears how to describe higher-lying states (c.f. Fig. 1). The natural choice in the framework of DFT seems to be TDDFT, as often applied in the literature but with mixed success. TDDFT is based on ground-state orbitals, where it can only describe single excitations if the usual energy-independent kernels are used[39]. The orbitals in the field of the core-hole may be better suited to describe the final state after photon absorption[16]. These orbitals already represent multiple excitations with respect to the ground-state orbitals. Along these lines, the maximum-overlap method[53] that considers the orbital changes under excitation is very successful, and sometimes even exceeds TDDFT accuracy in practical calculations.
-edge NEXAFS spectra have therefore been successfully calculated within a single-particle picture, where the corresponding orbitals are obtained within the field of an explicit core-hole. There are different approximations possible[17], and it has turned out that the transition-potential (TP) method[25, 26, 54] with and in Eq. (3) often leads to a rather good agreement with experiment[41, 13]. The excited core-hole (XCH) method, where and in Eq. (3) as appropriate for a neutral excitation, was also successfully applied[24, 55]. The system’s neutrality in XCH greatly benefits periodic calculations.
@@ III Computational settings
The electronic structure of the molecules and solids considered is described by DFT within the projector augmented wave method[56] as implemented in GPAW[57, 58, 59]. The exchange-correlation energy was evaluated as devised by Perdew, Burke and Ernzerhof[60].
@@
The smooth part of the Kohn-Sham orbitals as well as of the electron density are represented on real-space grids for most of the calculations. The grid spacing of the Kohn-Sham orbitals is 0.2 Å, while it is 0.1 Å for the electron density. The Kohn-Sham orbitals in the 2D system were represented by plane waves with an energy cut-off of 600 eV.
@@
We use the frozen-core approximation for orbitals not participating to the valence band, which are calculated for free atoms. The frozen-core approximation allows to describe the presence of a core hole obtained by a self-consistent free-atom calculation with a reduced occupation of the core state[26]. This approach explicitly considers a localized core hole and thus avoids problems with delocalized core-hole states in symmetric environments[17]. Details about the calculation of the shifts [Eq. (5)] are found in SI Sec. III.
@@
We contrast these single-particle calculations with linear-response TDDFT (LrTDDFT) calculations[35], where we include the relevant core-states into the valence orbitals. There is thus no explicit core-hole present in our LrTDDFT that is purely based on Kohn-Sham ground-state orbitals.
@@
The SO splittings are calculated through perturbation theory[61, 59] for free atoms. We have checked that the calculated SO-splittings vary very little compared to their values within the molecules, such that the assumption of a common SO splitting per element and core-state is a good approximation independent of the chemical environment (see SI Sec. IV for details).
@@
The oscillator strengths from Eq. (1) calculated via LrTDDFT or from Eq. (2) calculated via DFT use the corrected energies involving the shifts . In order to compare the spectra to experiment, these oscillator strengths are folded (convolved) by Cauchy distributions (Lorentzians) of variable width modeling to finite lifetime, vibrational contributions[62] and limited experimental resolution. The EEL spectra cover a larger range of excitation energies than the XAS spectra and higher-energy excitations are expected to have shorter lifetimes, resulting in larger broadening. This is taken into account by energy dependent broadening as detailed in SI. The convolution procedure generally results in intensities that are folded oscillator strengths (FOS) with the unit 1/eV.
@@ IV.1 Molecules
We generally evaluate the agreement between two spectra by calculating the coefficient of determination R2 value[65, 66, 67], where a value of unity describes perfect agreement. We allow for a variable width of the Cauchy distributions as well as for a variation in absolute energy by a shift to optimize agreement with experiment (see SI Sec. VI for details). Note that the shift is for a specific experiment, while the shift from Eq. (5) corrects the calculation to the average of high energy resolution experiments. We therefore regard our calculated and corrected energy as an accurate energy scale to which the specific experiment will be corrected.
@@
The better resolved spectrum of Stener et al.[64] shows a very small peak, about 2 eV lower than the main feature and well separated contributions above 464 eV. These features are well described by the lrTDDFT calculations resulting in a rather large R2. Similar agreement has been found from related calculations[64]. The energy of the first transition that is shifted to the corrected KS energy is indicated in the figure. This energy notably corresponds to a transition with very small oscillator strength, which therefore does not appear in the spectrum.
The single-particle XCH spectrum looks very different, however, and also shows a low R2. There are two strong peaks in the low-energy region caused by the contribution. These correspond to the unoccupied Ti 3 shell that is split due to the symmetry of the four Cl atoms. The corresponding 2 to 3 transitions are obviously coupled by LrTDDFT, but this is not the case within the single-particle picture. This coupling of empty states is known from atomic physics and is dubbed the multiplet effect in the literature[69].

### v2

@@ Abstract
Single-particle methods based on Kohn-Sham unoccupied states to describe near-edge X-ray absorption (XAS) spectra are routinely applied for the description of -edge spectra, as there is no complication due to spin-orbit (SO) coupling. - and -edge spectra are often addressed via variants of time-dependent density functional theory (TDDFT) based on SO calculations. Here, we present a computationally efficient implementation based on single-particle calculations with core holes within the frozen-core approximation. Combined with a semiempirical energy shift and a fixed spin-orbit splitting for each core level, this allows for a computationally cheap, while overall accurate prediction of experimental spectra on the absolute energy scale. The spectra are compared to about 40 times slower linear-response TDDFT calculations for molecules and show similar or even better match with experiment. An exception are multiplet effects that we analyze in detail and show that they cannot be covered by a single-particle approximation. A similar picture emerges for solids, where good qualitative and sometimes even quantitative agreement to experimental XAS and electron energy-loss spectra is achieved.
@@ I Introduction
There are several possible approaches for the simulation of XAS spectra as summarized in comprehensive reviews from different viewpoints[14, 15, 16, 17, 4]. The most accurate calculations are based on equation of motion-coupled cluster theory (EOM-CCSD), providing reliable spectra both in terms of shape[18, 19] as well as excitation energies[20, 21] despite of problems with orbital relaxation errors[22, 23]. However, such approaches are only applicable to smaller molecules due to their extreme computational cost and the resulting scalability limitations.
The workhorse for describing excitations in larger systems is density functional theory (DFT). Here, the involvement of (highly) excited states suggests the use of time-dependent DFT (TDDFT)[24, 25]. However, TDDFT is also computationally expensive, and it has limitations in particular for the description of solids in combination with local- or semi-local functionals[26]. An alternative to TDDFT are single-particle core-hole methods[17, 27, 28] where the unoccupied states in the field of a core hole are interpreted as excitations. These methods are usually applied to -edge spectra[28, 29, 30, 13, 17, 30] due to the difficulties of dealing with the angular momentum of the initial state and with spin-orbit (SO) coupling[17], although several -edge EELS simulations using the single-particle density of states[31] have notably been reported[32, 33, 34, 11].
@@
In this article, we apply single-particle methods to simulate - and -edge NEXAFS spectra, specifically focusing on 2 and 3 transitions. We compare our approach to experiment for molecules as well as for solids in order to explore their strengths and weaknesses, and also compare to TDDFT spectra in the case of molecules. The manuscript first discusses the picture underlying the description of NEXAFS spectra with an emphasis on comparing the transition from the many-particle to the single-particle viewpoint. A simple, but effective treatment of SO coupling as a constant shift is also presented. Then the predictions of the methods are compared to experiment both for molecules as well as for solid systems. After showing that our methodology works extremely well also for single-atom EEL spectra, we finish with conclusions.
@@ II X-ray absorption
The process of X-ray absorption can be interpreted in different approximations, as depicted schematically in Fig. 1. Generally, an X-ray photon is completely absorbed by the molecule or solid under observation, leading to a highly excited state of the system due to the large photon energy.
We first discuss the process within the most accurate many-particle picture in terms of many particle ground and excited states. The absorption cross-sections are proportional to the unitless oscillator strength[27, 38, 39, 40]
where is the electron mass, the unit charge, the reduced Planck constant and are Cartesian coordinates. is the energy of the ground state, the energy of the core-excited state of index (there are many excited states forming the full XAS-spectrum) and denote their corresponding many-electron wave functions. The dipole operator in length form contains the contributions of all electrons. The relevant coordinate is chosen by the polarization of the X-ray photon . A specific appears in angle-dependent X-ray spectroscopy[41] or X-ray circular dichroism[16]. A similar expression involving many-electron wave functions appears in the GW approximation[42] resulting in Eq. (1) for vanishing momentum transfer by the absorbed photon. The target molecules or solids are usually in random orientations, such that the averaged intensity, i.e. is measured.
The calculation of transition energies and oscillator strengths in Eq. (1) requires many body energies and wave functions. Many-electron methods are very demanding requiring expansion in a large number of single particle orbitals or many particle-hole pairs. Most exact, but also most demanding is the use of many-body wave functions as provided by e.g. coupled cluster (CC) descriptions for ground and excited states. The equivalent could be achieved by TDDFT in case the exact (frequency dependent[43]) exchange-correlation kernel would be known. Practical TDDFT calculations, in particular based on gradient-corrected kernels as applied here, do not, however, necessarily lead to accurate transition energies[44].
A huge simplification is achieved when XAS is interpreted within the single-particle picture as transitions of a core electron to an empty orbital (or the conduction band for solids). The oscillator strength reduces in this picture to[45]
@@
Eq. (2) is a harsh approximation as the energies and wave-functions of many-electron excitations are replaced by the energy and orbital of a single, seemingly independent unoccupied state, that has to describe all excitations to itself irrespective where the excitation came from. We will see below, that this approximation indeed breaks down for strongly correlated systems[46].
These single-particle energies may be again evaluated in different approximations that mainly differ by the occupation of the core state as well as by the occupation of the lowest unoccupied state (LUMO). These choices not only influence the energy values, but also change the single-particle states in Eq. (2), where we have suppressed the labels for clarity. More details about the implementation can be found in the Supplementary Material (SM) Sec. II.
Regardless of whether single-particle or many-particle pictures are adopted, the calculated spectra are usually rigidly shifted to match their lowest energy peak with that of the experiment[47]. With the exception of some reported computationally expensive EOM-CCSD calculations[24] these shifts often amount to several eV.
@@
that can be used to correct the raw calculated energies as indicated in Fig. 1. This method has successfully predicted the XPS spectra of ionic clusters [51] and carbon materials[52, 53], as well as to S 1 NEXAFS spectra [13, 3]. We will see below that there is also substantial uncertainty in experimental absolute energies. The shift can therefore be seen as an average over experimental knowledge. The shift depends on core state and depends on the exchange-correlation functional. Optimizing the shift to experiment therefore also corrects for inaccuracies of the functional. There are hints that at least a large part of the shift stems from the use of the frozen-core approximation, as all-electron calculations seem to have negligible shifts at least for light elements[54, 55].
@@
Having obtained an accurate value for the first transition energy, i.e. to the LUMO in NEXAFS spectra, the question appears how to describe higher-lying states (c.f. Fig. 1). The natural choice in the framework of DFT seems to be TDDFT, as often applied in the literature but with mixed success. TDDFT is usually based on ground-state orbitals, where it can only describe single excitations if the usual energy-independent kernels are used[43]. The orbitals in the field of the core-hole may be better suited to describe the final state after photon absorption[16]. These orbitals already represent multiple excitations with respect to the ground-state orbitals and have been demonstrated to improve TDDFT XAS calculations[58]. Along these lines, methods that consider orbital changes under excitation[59, 60] are very successful and sometimes even exceed standard TDDFT accuracy in practical calculations.
-edge NEXAFS spectra have therefore been successfully calculated within a single-particle picture, where the corresponding orbitals are obtained within the field of an explicit core-hole. There are different approximations possible[17], and it has turned out that the transition-potential (TP) method[28, 29, 61] with and in Eq. (3) often leads to a rather good agreement with experiment[45, 13]. There are rationalizations for other values[62, 63] in extension of Slater’s original transition state concept. The latter is based on setting an occupation of in the excited orbital of interest, which is further approximated in the TP method to in order to create a single potential for all excitations[28]. In the viewpoint of the various approximations adopted, we generally use as in original TP, but treat it as free parameter in some calculations. The excited core-hole (XCH) method, where and in Eq. (3) as appropriate for a neutral excitation, was also successfully applied[27, 64]. The system’s neutrality in XCH greatly benefits periodic calculations.
@@ III Computational settings
The electronic structure of the molecules and solids considered is described by DFT within the projector augmented wave method[65] as implemented in GPAW[66, 67, 68]. The exchange-correlation energy was evaluated as devised by Perdew, Burke and Ernzerhof (PBE)[69]. PBE is a robust functional derived from basic principles which is applied for molecular systems as well as solids. We have good experience using PBE for the description of core-hole spectra on the absolute energy scale by applying a semi-empirical energy correction fitted for this functional[48, 51, 53, 13].
@@
GPAW offers the possibility to represent the smooth part of the Kohn-Sham orbitals on real-space grids, momentum space (plane wave) grids or linear combination of atomic orbitals. The electron density is represented on real-space grids, where we chose a grid spacing of 0.1 Å in our calculations. The Kohn-Sham orbitals are also represented on real-space grids with 0.2 Å grid-spacing. The Kohn-Sham orbitals in the 2D system were represented by plane waves with an energy cut-off of 600 eV. The use of grids in real or momentum space (plane waves) offers a facile and smooth convergence to the complete basis set limit[66, 70]. Equidistant grids also allow for very effective parallelization.[67]
We use the frozen-core approximation[66, 67] for orbitals not participating to the valence band, which are calculated for free atoms. This approximation allows to use coarser grids while still catching all orbitals necessary to cover chemical interactions. In explicit the frozen core states are 1s2=[He] for C and O, [He]2s22p6=[Ne] for Si, S, Cl and Ti, [Ne]3s2 for Ni, [Ne]3s23p63d10=[Ar]3d10 for Sr, and [Ar]3d104s24p6=[Kr] for Sn.
@@
The frozen-core approximation has the additional benefit in that it allows to describe the presence of a core hole obtained by a self-consistent free-atom calculation with a reduced occupation of the core state[29]. This approach explicitly considers a localized core hole and thus avoids problems with delocalized core-hole states in symmetric environments[17], despite that solutions to this problem exist[71, 72]. The price to pay is the absence of core-hole state relaxation in a chemical environment different to an isolated atom changing the energetics[54]. This is considered and corrected by our semiempirical shifts that were fitted for these fixed core-holes. Details about the calculation of the shifts [Eq. (5)] are found in SM Sec. III.
@@
We contrast these single-particle calculations with linear-response TDDFT (LrTDDFT) calculations[38], where we include the relevant core-states into the valence orbitals. There is thus no explicit core-hole present in our LrTDDFT that is purely based on Kohn-Sham ground-state orbitals. LrTDDFT in combination with gradient-corrected kernels as applied here though the PBE functional is known have severe problems in describing charge-transfer excitations.[73] We therefore contrast selected spectra calculated with a range-separated functional based on the Yukawa potential (LCY-PBE)[74, 75] which is known to perform better in this respect[75]. The Yukawa damping parameter is obtained by requiring the equality of the highest molecular orbital’s energy with the negative of the ionization potential[76, 77] as valid in exact DFT[78].
@@
The SO splittings for each element are calculated through perturbation theory[79, 68] for free atoms as detailed in SM. We have checked that the calculated SO-splittings vary very little compared to their values within the molecules, such that the assumption of a common SO splitting per element and core-state is a good approximation independent of the chemical environment (see SM Sec. IV for details). The oscillator strengths are calculated without spin-orbit effects and the full spectrum is constructed assuming the same spectrum for the two contributions separated in energy by the SO splitting and scaled by the corresponding weight as given in Tab. S1 in SM.
@@
The oscillator strengths from Eq. (1) calculated via LrTDDFT or from Eq. (2) calculated via DFT use the corrected energies involving the shifts . In order to compare the spectra to experiment, these oscillator strengths are folded (convolved) by Cauchy distributions (Lorentzians) of variable width modeling to finite lifetime, vibrational contributions[80] and limited experimental resolution. The EEL spectra cover a larger range of excitation energies than the XAS spectra and higher-energy excitations are expected to have shorter lifetimes, resulting in larger broadening. This is taken into account by energy dependent broadening as detailed in SM. The convolution procedure generally results in intensities that are folded oscillator strengths (FOS) with the unit 1/eV.
@@ IV.1 Molecules
We generally evaluate the agreement between two spectra by calculating the coefficient of determination R2 value[83, 84, 85], where a value of unity describes perfect agreement. We allow for a variable width of the Cauchy distributions as well as for a variation in absolute energy by a shift to optimize agreement with experiment (see SM Sec. VI for details). Note that the shift is for a specific experiment, while the shift from Eq. (5) corrects the calculation to the average of high energy resolution experiments. We therefore regard our calculated and corrected energy as an accurate energy scale to which the specific experiment will be corrected.
@@
The better resolved spectrum of Stener et al.[82] shows a very small peak, about 2 eV lower than the main feature and well separated contributions above 464 eV. These features are well described by the lrTDDFT calculations resulting in a rather large R2. Similar agreement has been found from related calculations[82]. The energies of the first transitions that are shifted to the corrected KS energy for and to KS plus the SO splitting for are indicated in the figure. This energy notably corresponds to a transition with very small oscillator strength, which therefore does not appear in the spectrum. We note, that the shift eV to align the experimental spectrum with the LrTDDFT spectrum is smaller than the corresponding alignment between the two experiments.
The single-particle XCH spectrum looks very different, however, and also shows a low R2. The TP spectrum is very similar and can be found in SM. There are two strong peaks in the low-energy region caused by the contribution. These correspond to the unoccupied Ti 3 shell that is split due to the symmetry of the four Cl atoms. The corresponding 2 to 3 transitions are obviously coupled by LrTDDFT, but this is not the case within the single-particle picture. This coupling of empty states is known from atomic physics and is dubbed the multiplet effect in the literature[87].
In order to understand the origin of the multiplet effect, we compare the 2p to 3d transitions to the 3s to 3d transitions in Fig. 4. The density of excited states (DOS) corresponding to 3s to 3d transitions show two peaks due to the splitting of the d-states in the tetrahedral symmetry of the Cl atoms in TiCl4. There is some small transition probability for one type of these excitations due to mixing of p-states in this otherwise dipole-forbidden transition. This is reflected both in the single particle transitions (KS) as well as in the excitations from LrTDDFT. I.e. TDDFT merely leads to a slight energy shift here.
The situation is remarkably different when comparing KS and LrTDDFT for 2p to 3d transitions in Fig. 4 e-h). We restrict to the main contribution 2p3/2 contribution, the part is found at higher energies. The KS single-electron picture shows two transition energies split like the 3s to 3d transitions as all relevant 2p states have the same energy. Now both DOS peaks have non-vanishing oscillator strength in KS giving the two prominent peaks seen in Fig. 2b). The DOS of LrTDDFT is very different in that more peaks are built from the 15 2p to 3d transitions. The most dramatic effect is seen for the oscillator strength that now is practically only finite for the highest energy transitions, leading to one strong main peak as seen in Fig. 2a) and in experiment.
We note, that this effect cannot be described within an independent orbital picture. TiCl4 therefore represents a strongly correlated system in the empty states. The excitations in TiCl4 change according to "where they came from", such that there are no independent -orbitals that are able to describe all transitions irrespective from which orbital they are filled.
@@
The broadening in the LrTDDFT spectrum is caused by the gradient-corrected functional PBE we are using as is demonstrated by a comparison applying the range separated LCY-PBE functional in Fig. 3d). The LCY-PBE spectrum is more peaked and at least partly resembles the experimental peak structure from experiment. We note, that the use of Hartree-Fock exchange as required in LCY-PBE is extremely expensive in grid based approaches due to the need for an explicit calculation of the exchange integrals. These can be calculated partly analytically in basis set calculations, which is not possible here. The LCY-PBE LrTDDFT calculation used 260 CPU days, while the PBE LrTDDFT calculation finished in 9 CPU hours, i.e. three orders of magnitude faster. The TP and XCH calculations finished in roughly 14 CPU minutes, giving a speed-up factor of 40 relative to PBE LrTDDFT and 30.000 relative to LCY-PBE and demonstrating the cost-effectiveness of the single-particle methods.